Psychology 406 Computer Assignment #2

Use your old data set from Computer Assignment #1 in (a), (b), and (c):

a) Begin by listing the contents of your SYSTAT file for the benefit of the grader.

b) From the STATISTICS menu, use the T-TEST command to carry out two paired t-tests: for the pennies versus dimes within room 1 and for pennies versus dimes within room 2 (thus, treat pennies and dimes as if matched by case).

c) From the STATISTICS menu, use the T-TEST command to carry out two two-independent sample t-tests: use pennies for room 1 versus room 2 and dimes for room 1 versus room 2 (thus, rooms are treated as independent samples).

The data to be used in (d) through (h) are on 170 entering freshman at a large public university, and is available from the web site:

http://cda.psych.uiuc.edu/psych 406 material/P406data.syd

The data set involves five variables: self-reported SAT verbal score (SATV), self-reported SAT math score (SATM), actual SAT verbal score (RSATV), sex (SEX), and high school grade point average (HSGPA). The three SAT scores are given as two digits (the usual 0 could be added at the end of each), sex involves a “one” sex and a “zero” sex, and HSGPA is given as one of nine levels (from 0 to 8). You will need to define for yourself a variable describing the difference between the actual SAT verbal score and the reported SAT verbal score (SATVDIFF).

d) Using the STATISTICS menu, obtain all the usual descriptive statistics for your data set — include SATVDIFF.

e) Using the CORRELATIONS menu, obtain the simple Pearson productmoment correlations between all the variables (including SATVDIFF).

f) Using the REGRESSION menu, carry out simple regression analyses between SATVDIFF and each of the variables SATM, SEX, and HSGPA. Interpret.

g) Carry out a multiple regression analysis (using REGRESSION) predicting SATVDIFF from SATM, SEX, and HSGPA. Save the residuals and list these along with leverages, studentized residuals, Cook’s D values, and the standard error of the predicted values. Interpret.

h) Carry out a stepwise regression (using REGRESSION) in predicting SATVDIFF from SATM, SEX, and HSGPA, and interpret.